OpenStudy (anonymous):

If f(x) varies directly with x2, and f(x) = 72 when x = 6, find the value of f(3) Help?

OpenStudy (adamaero):

f(3) means when x = 3

OpenStudy (adamaero):

72 = 2x

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

Would it be 18

OpenStudy (anonymous):

@adamaero

OpenStudy (adamaero):

for x = 3 or x = 6? but actually I think I read the problem too fast, I lead you down the wrong path... let me check...

OpenStudy (anonymous):

x=6. I think.

OpenStudy (anonymous):

I'm just generally confused about this entire thing lol.

OpenStudy (adamaero):

are you sure it's not x^2 ?

OpenStudy (anonymous):

Oh, sh** lol. I forgot to put the little ^. Yes, it's x^2. Sorry about that.. haha

OpenStudy (adamaero):

lol, you can't swear on here anyways flutter that feather

OpenStudy (anonymous):

I know haha, I've tried it before.

OpenStudy (adamaero):

wio's already on the way answering it for you: http://openstudy.com/study#/updates/54877db4e4b0c0f1a1527fc0

OpenStudy (adamaero):

here is someone else answering it for you: http://openstudy.com/study#/updates/54877e6be4b0c0f1a1528099

OpenStudy (anonymous):

Oh. Okay. Thanks so much lol

myininaya (myininaya):

@Chelleyhi are you there?

myininaya (myininaya):

\[f(x)=kx^2 \] this is the first translation you should get

myininaya (myininaya):

now we can find the constant k given the second sentence the second sentence says we have f(x)=72 when x=6 so we have 72=k(6)^2

myininaya (myininaya):

solving that for k gives us what?

OpenStudy (anonymous):

18?

myininaya (myininaya):

72/36= is that 18

myininaya (myininaya):

let me think 72/36=2

myininaya (myininaya):

isn't it?

myininaya (myininaya):

since 36+36=72 there are 2(36)=72 so 2=72/36

myininaya (myininaya):

\[f(x)=kx^2 \\ \text{ so k is 2} \\ f(x)=2x^2 \]

myininaya (myininaya):

now just plug in 3 to find f(3)

myininaya (myininaya):

f(3)=2(3)^2

myininaya (myininaya):

oh do you mean f(3)=18 if so yes

myininaya (myininaya):

since 2*3^2=2*9=18

OpenStudy (anonymous):

Thank you for your help, I did a bunch of googling and finally figured out how to do it.

myininaya (myininaya):

Good job you! :)